If 115 is multiplied by certain number,it is increased by2760.find the number 

Grade:6

3 Answers

Varun
6 years ago
Solution :115x =x+2760 ,,114x=2760,,,,x=2760/114 I HOPE ITS CLEAR .ALL THE BEST.....................,..................................
Shahid
6 years ago
Dear Arnab , 
Let The Number be “x”
Then According to the Question... 
(115) (x) = (x) + 2760 …
114x = 2760 
x = 24.2105
x \simeq 24.2
 
Hope It Helped...
Thank You!!! 
Madhurima Majee
2 years ago
Hi.....
Let the number is x
Then According to the Question... 
115x = (2760+115)
   x = 2875/115 =25.           
 
Hope It Helped...
Thank You!!!